I know I'm solving for x, so I think I set up the equation like
(2x+5)+(3x+1)+(13)+(5x-7)=? My problem is that I don't know what I set it to equal to.
Photobucket
: I know I'm solving for x, so I think I set up the equation like
(2x+5)+(3x+1)+(13)+(5x-7)=? My problem is that I don't know what I set it to equal to.
Photobucket

Answer by jim_thompson5910(9929)

Since the diagonals of a rectangle are equal, this means that we have the equation

(2x+5)+(5x-7)=(3x+1)+(13)


2x+5+5x-7=3x+1+13 Remove the parenthesis



7x-2=3x+1+13 Combine like terms on the left side


7x-2=3x+14 Combine like terms on the right side


7x=3x+14+2Add 2 to both sides


7x-3x=14+2 Subtract 3x from both sides


4x=14+2 Combine like terms on the left side


4x=16 Combine like terms on the right side


x=(16)/(4) Divide both sides by 4 to isolate x



x=4 Divide

--------------------------------------------------------------
Answer:
So our answer is x=4